4. The body with mass 3kg, is located in the sloping plane that forms the angle 45 k, with the horizontal direction. What must be the force F, exerted on the body, for it to remain in balance. The coefficient of friction of the body rest with the plane is 0.2


1
Expert's answer
2021-03-26T11:26:31-0400
"F=mg\sin{45}-\mu mg \cos{45}\\
F=mg(\sin{45}-\mu \cos{45})\\F=(3)(9.8)(\sin{45}-0.2\cos{45})=16.6\ N"